# WASUP: Interpretable Classification with Weight-Input Alignment and Class-Discriminative SUPports Vectors

### v1

@@ 5 Discussion and Conclusion
Current inherently interpretable models either lack global explanations, fail to explain the internal mechanisms of the deep learning model, or cannot provide explanations that faithfully explain their decision-making. To this end, we proposed WASUP for interpretable classification by providing faithful local and global explanations by leveraging case-based reasoning and weight-input alignment achieved with B-cos neural networks. Our theoretical analysis shows that our proposed method’s explanations satisfy the required axioms for faithfulness. We evaluated WASUP on three tasks using three backbone architectures, demonstrating how the complex mechanisms learned by neural networks can be decomposed into intuitive explanations with limited complexity.
Additionally, we exemplified how its global and local explanation properties enable effective model debugging from a developer perspective while giving intuitive explanations of its decision-making. We focused on creating a model that yields comprehensible explanations, which can, if desired, decompose each test contribution into individual components based on the similarity measure. This generates additional contribution maps for each support sample, facilitating further validation in cases with high intra-class variability, albeit with an increase in complexity.
While k-means clustering is fast to optimize, WASUP introduces additional overhead by requiring the computation of support vectors, which is a potential limitation in cases with many classes. However, as demonstrated with the Stanford Dogs dataset comprising 120 classes, WASUP is readily applicable out of the box to datasets with hundreds of classes. Moreover, note that the number of classes in decision-critical fields such as medicine is generally smaller. In future work, we are curious to extend WASUP by conformal prediction to automatically show the explanations of a set of possible classes as opposed to the most likely one in single-label tasks.
@@ A.1 On the Faithfulness of B-cos Explanations
B-cos modles are piece-wise linear models that allows to summarize each input as an input-dependent linear transform $f({x})=\mathbf{W}_{1\rightarrow L}(x)x$, derived in Eq. 3. Explanations are computed in terms of the weight matrix $\mathbf{W}_{1\rightarrow L}(x)$ by multiplying each feature with its matrix weight $\phi(x)_{i}=[\mathbf{W}_{1\rightarrow L}]^{T}_{j}\odot x]_{(ch,i,j)}$. To show that computing explanations for a forward pass is faithful (i.e., satisfies below six axioms introduced by Sundararajan et al. (2017)), we need to show that they hold considering the transformation matrix $\mathbf{W}_{1\rightarrow L}(x)$. While $\mathbf{W}_{1\rightarrow L}(x)$ is indeed input-dependent, it is effectively fixed during the computation of the explanation for its specific input $x$. This is because the explanations $\phi(f,X)$ aim to attribute the output $f(x)$ to the input features of $x$ using the weights at that point. Therefore, we assume a fixed $\mathbf{W}_{1\rightarrow L}(x)$ when explaining the forward pass of an input $x$, and acknowledge that while $\mathbf{W}_{1\rightarrow L}(x)$ varies across different inputs, it remains constant within the context of computing local explanations $\phi(f,x)$ for a particular $x$. In addition, we consider the case of single-class prediction, where the weight matrix simplifies to a single-column form ($\mathbf{W}_{1\rightarrow L}(X)\in\mathbb{R}$) for in input vector $X\in\mathbb{R}$. This allows us to focus on the core properties of the explanation mechanism without loss of generality since explaining the prediction of a class in the multiclass setting is analogous to explaining its corresponding column vector of the matrix $\mathbf{W}_{1\rightarrow L}(x)$.
We abbreviate $\mathbf{W}_{1\rightarrow L}(x)$ with $W(X)$ for ease of notation in the following. Let  WASUP be defined as $f(X)=W(X)\cdot X$ and its contributions similarily by $\phi(f,X)=W(X)\odot X=(W(X)_{1}\times X_{1},\dots,W(X)_{N}\times X_{N})$ (remark that $\cdot$ denotes the scalar product and $\odot$ the element-wise multiplication, $\times$ the multiplication of real numbers).
@@ A.2 On the Faithfulness of support vectors
Justification for ReLU in WASUP
In contrast to other functions that could be used to implement $\oplus$, the ReLU activation does not alter the range of the input to potentially huge or tiny numbers like, e.g. the exponential. However, ReLU activations are commonly scrutinized for their role in setting negative activations to zero, which can cause deep neural networks to interpret this suppression as the absence of a feature. Thus, many XAI methods (e.g. gradient-based) yield misleading attribution maps, as they cannot account for this type of contribution. However, this phenomenon is tightly controlled in WASUP, where the sole operation following the ReLU activation is a similarity computation between the latent vectors of two images. As a result, the absence of a feature does not, by design, affect the probability of class logits, rendering it valid to disregard any input and its explanation from neurons corresponding to such features.
@@ A.3 Implementation and Training Details
All models were optimized using the AdamW optimizer with fixed hyperparameters: weight decay set to 0.0, betas configured to (0.9, 0.999), and an epsilon value of 1e-08. We conducted an extensive grid search over the hyperparameter space (see Tab. A.2) to identify the optimal settings based on accuracy. The training employed a custom learning rate scheduling policy, which began with a linear warm-up phase over the first two epochs, increasing the learning rate from 10% to the initial learning rate. This was followed by maintaining the initial learning rate until 50% of the total training iterations were completed. Subsequently, the learning rate was decayed by a factor of 0.5 at every subsequent 10% milestone of the total iterations. Batch sizes were set to 32 for all datasets except RSNA, which utilized a batch size of 64. The number of training epochs was standardized to 50 across all datasets, with RSNA trained for 100 epochs. Pretrained feature extractor weights were sourced from torchvision https://pytorch.org/vision/stable/index.html and the official B-cos repository https://github.com/B-cos/B-cos-v2 for all experiments except those involving RSNA.
@@
Model-specific parameters included setting B-cos $B=2$ and using three support samples. For WASUP, the grid search explored learning rates of 0.01, 0.003, 0.001, and 0.0003 for the RSNA dataset, and 0.0022, 0.001, 0.00022, 0.0001, 0.000022, and 0.00001 for other datasets. Temperature parameters tested were 10, 30, and 50 for RSNA, and 10 and 30 for the remaining datasets. For BlackBox models, the grid search included the same learning rates as WASUP and additionally varied the weight decay parameter, evaluating values of 0.0 and 0.0001. The best-performing models were selected based on their accuracy metrics.
Table A.2: Best configurations and performance of each model. WASUP consistently performs marginally better ($\geq+0.29\%$) than the black box with the same backbone on Pascal VOC. On Stanford Dogs, the black box outperforms WASUP by $3.7\%$. On RSNA, the difference ($0.65\%$) marginally favors the black box, although a repeatability study with different random weight initializations is needed to confirm this finding (note that models were only randomly initialized in the RSNA experiment).
Method Model Best LR Best WD Best Temp. Acc (%)
BlackBox VOC - DenseNet121 $2.2\times 10^{-5}$ 0.0 - 96.42
VOC - ResNet50 $1.0\times 10^{-5}$ 0.0 - 96.71
VOC - VitC $1.0\times 10^{-4}$ 0.0 - 96.35
StanfordDogs - ResNet50 $1.0\times 10^{-5}$ 0.0 - 87.16
RSNA - DenseNet121 $3.0\times 10^{-4}$ $1.0\times 10^{-4}$ 79.78
WASUP VOC - DenseNet121 $2.2\times 10^{-4}$ 30 96.73
VOC - ResNet50 $1.0\times 10^{-4}$ 30 97.00
VOC - VitC $2.2\times 10^{-5}$ 10 96.72
StanfordDogs - ResNet50 $1.0\times 10^{-4}$ 10 83.46
RSNA - DenseNet121 $3.0\times 10^{-4}$ 50 79.13
@@
Table A.3: Dataset statistics of the RSNA dataset.
Dataset Split Patient Sex Target Number of Samples
Training Set F No Opacity 6,770
@@ A.4.4 RSNA - DenseNet
Figure A.13: Prototypes of WASUP trained with a DenseNet121 backbone on RSNA. Bounding boxes indicate where medical staff found occlusion of the lungs. The first three columns are the support samples of controls, the next three are support samples with occlusion present.
@@
A.4.5 VOC -DenseNet121 Prototypes and B-cos Similarities
Figure A.15: Global explanations for WASUP trained on Pascal VOC with a DenseNet121 backbone. The respective second rows show the raw contribution maps $\phi^{L}_{j=c}$, in which red denotes positive contributions and blue (absent) negative contributions. The respective third columns present RGBA explanations.
Figure A.16: Heatmap of Intra-class and Inter-class B-cos Similarities for WASUP trained on Pascal VOC with a DenseNet121 backbone.

### v2

@@ 5 Discussion and Conclusion
Current inherently interpretable models either lack global explanations, fail to explain the internal mechanisms of the deep learning model, or cannot provide explanations that faithfully explain their decision-making. To this end, we proposed SIC for similarity-based interpretable classification by providing faithful local, global, and pixel-level explanations. Our theoretical analysis proved that our proposed method’s explanations satisfy established axioms for explanations, namely  CompletenessSensitivityImpementation InvarianceDummyLinearity, and  Symmetry-Preserving.
We evaluated our method, SIC, across three tasks using three distinct backbone architectures, demonstrating that the complex mechanisms learned by neural networks can be decomposed into intuitive, low-complexity explanations. Our practical evaluation employing the FunnyBirds framework further illustrated that the careful design of SIC manifests in its explanations. Notably, our method outperformed established inherently interpretable approaches on the majority of evaluation metrics, while its classification accuracy remained comparable to that of black-box models in two out of three real-world datasets.
We demonstrated that its intuitive global and local explanations not only facilitate effective model debugging for developers but also simplify the understanding of its decision-making process for end-users.
While k-means clustering is fast to optimize, SIC introduces additional overhead by requiring the computation of support vectors, which is a potential limitation in cases with many classes. However, as demonstrated with the Stanford Dogs dataset comprising 120 classes, SIC is readily applicable to datasets with hundreds of classes. Moreover, we note that the number of classes in decision-critical fields such as medicine is generally smaller.
In summary, our results demonstrated that embedding interpretability within complex deep neural networks not only achieves competitive performance on challenging benchmarks but also paves the way for more transparent and trustworthy computer vision systems.
@@ A.1 On the Faithfulness of B-cos Explanations
B-cos modles are piece-wise linear models that allows to summarize each input as an input-dependent linear transform $f({x})=\mathbf{W}_{1\rightarrow L}(x)x$, derived in Eq. 2. Explanations are computed in terms of the weight matrix $\mathbf{W}_{1\rightarrow L}(x)$ by multiplying each feature with its matrix weight $\phi(x)_{i}=[\mathbf{W}_{1\rightarrow L}]^{T}_{j}\odot x]_{(ch,i,j)}$. To show that computing explanations for a forward pass is faithful (i.e., satisfies below six axioms introduced by Sundararajan et al. [30]), we need to show that they hold considering the transformation matrix $\mathbf{W}_{1\rightarrow L}(x)$. While $\mathbf{W}_{1\rightarrow L}(x)$ is indeed input-dependent, it is effectively fixed during the computation of the explanation for its specific input $x$. This is because the explanations $\phi(f,X)$ aim to attribute the output $f(x)$ to the input features of $x$ using the weights at that point. Therefore, we assume a fixed $\mathbf{W}_{1\rightarrow L}(x)$ when explaining the forward pass of an input $x$, and acknowledge that while $\mathbf{W}_{1\rightarrow L}(x)$ varies across different inputs, it remains constant within the context of computing local explanations $\phi(f,x)$ for a particular $x$. In addition, we consider the case of single-class prediction, where the weight matrix simplifies to a single-column form ($\mathbf{W}_{1\rightarrow L}(X)\in\mathbb{R}$) for in input vector $X\in\mathbb{R}$. This allows us to focus on the core properties of the explanation mechanism without loss of generality since explaining the prediction of a class in the multiclass setting is analogous to explaining its corresponding column vector of the matrix $\mathbf{W}_{1\rightarrow L}(x)$.
We abbreviate $\mathbf{W}_{1\rightarrow L}(x)$ with $W(X)$ for ease of notation in the following. Let  SIC be defined as $f(X)=W(X)\cdot X$ and its contributions similarily by $\phi(f,X)=W(X)\odot X=(W(X)_{1}\times X_{1},\dots,W(X)_{N}\times X_{N})$ (remark that $\cdot$ denotes the scalar product and $\odot$ the element-wise multiplication, $\times$ the multiplication of real numbers).
@@ A.2 On the Faithfulness of support vectors
Justification for ReLU in SIC
In contrast to other functions that could be used to implement $\oplus$, the ReLU activation does not alter the range of the input to potentially huge or tiny numbers like, e.g. the exponential. However, ReLU activations are commonly scrutinized for their role in setting negative activations to zero, which can cause deep neural networks to interpret this suppression as the absence of a feature. Thus, many XAI methods (e.g. gradient-based) yield misleading attribution maps, as they cannot account for this type of contribution. However, this phenomenon is tightly controlled in SIC, where the sole operation following the ReLU activation is a similarity computation between the latent vectors of two images. As a result, the absence of a feature does not, by design, affect the probability of class logits, rendering it valid to disregard any input and its explanation from neurons corresponding to such features.
@@ A.3 Implementation and Training Details
All models were optimized using the AdamW optimizer with fixed hyperparameters: weight decay set to 0.0, betas configured to (0.9, 0.999), and an epsilon value of 1e-08. For FunnyBirds, we selected the default parameters of the framework: a learning rate of 0.001, and weight decay of 0.0, trained for 100 epochs with a temperature of 30 and a batch size of 8. For all other experiments, we conducted an extensive grid search over the hyperparameter space (see Tab. A.2) to identify the optimal settings based on accuracy. The training employed a custom learning rate scheduling policy, which began with a linear warm-up phase over the first two epochs, increasing the learning rate from 10% to the initial learning rate. This was followed by maintaining the initial learning rate until 50% of the total training iterations were completed. Subsequently, the learning rate was decayed by a factor of 0.5 at every subsequent 10% milestone of the total iterations. Batch sizes were set to 32 for all datasets except RSNA, which utilized a batch size of 64. The number of training epochs was standardized to 50 across all datasets, with RSNA trained for 100 epochs. Pretrained feature extractor weights were sourced from torchvision https://pytorch.org/vision/stable/index.html and the official B-cos repository https://github.com/B-cos/B-cos-v2 for all experiments except those involving RSNA. We train ProtoPNet with the default training recipe provided on the official GitHub implementation (https://github.com/cfchen-duke/ProtoPNet) and adapted the code for VOC for multi-label classification. Note that the NW-Head is not applicable to multi-label classification tasks as it relies on softmax for normalization, which can not be adapted with minor modifications as done with ProtoPNet.
@@
Model-specific parameters included setting B-cos $B=2$ and using three support samples. We train the NW-Head in the replace cluster mode. For SIC, the grid search explored learning rates of 0.01, 0.003, 0.001, and 0.0003 for the RSNA dataset, and 0.0022, 0.001, 0.00022, 0.0001, 0.000022, and 0.00001 for other datasets. Temperature parameters tested were 10, 30, and 50 for RSNA, and 10 and 30 for the remaining datasets. For BlackBox models, the grid search included the same learning rates as SIC and additionally varied the weight decay parameter, evaluating values of 0.0 and 0.0001. The best-performing models were selected based on their accuracy metrics.
Method Dataset Backbone Best LR Best WD Best Temp. Acc (%)
BlackBox VOC DenseNet121 $2.2\times 10^{-5}$ 0.0 - 96.42
VOC ResNet50 $1.0\times 10^{-5}$ 0.0 - 96.71
VOC VitC $1.0\times 10^{-4}$ 0.0 - 96.35
StanfordDogs ResNet50 $1.0\times 10^{-5}$ 0.0 - 87.16
RSNA DenseNet121 $3.0\times 10^{-4}$ $1.0\times 10^{-4}$ 79.78
NW-Head [31] StanfordDogs ResNet50 $1.0\times 10^{-5}$ $0.0$ 85.16
RSNA DenseNet121 $3.0\times 10^{-4}$ $1.0\times 10^{-4}$ 79.30
ProtoPNet [8] VOC DenseNet121 - - - 91.19
VOC ResNet50 - - - 89.31
StanfordDogs ResNet50 - - - 49.24
RSNA DenseNet121 - - - 78.68
BagNet [6] VOC BagNet17 $2.2\times 10^{-5}$ $1.0\times 10^{-4}$ 94.56
StanfordDogs BagNet17 $2.2\times 10^{-5}$ 0.0 - 62.32
RSNA BagNet17 $3.0\times 10^{-4}$ $1.0\times 10^{-4}$ 77.70
B-cos [5] VOC DenseNet121 $1.0\times 10^{-4}$ 96.48
VOC ResNet50 $1.0\times 10^{-4}$ 96.85
VOC VitC $2.2\times 10^{-5}$ 96.79
StanfordDogs ResNet50 $1.0\times 10^{-4}$ 85.54
RSNA DenseNet121 $1.0\times 10^{-2}$ 78.73
SIC VOC DenseNet121 $2.2\times 10^{-4}$ 30 96.73
VOC ResNet50 $1.0\times 10^{-4}$ 30 97.00
VOC VitC $2.2\times 10^{-5}$ 10 96.72
StanfordDogs ResNet50 $1.0\times 10^{-4}$ 10 83.46
RSNA DenseNet121 $3.0\times 10^{-4}$ 50 79.13
Table A.2: Best configurations and performance of each model.
@@
Split Sex Target Num. Samples
Train F No Opacity 6,770
@@ F Opacity 629
We follow the FunnyBirds framework and use the default data processing pipeline and compute the metrics with the unsmoothed explanations $\phi$, i.e. the alpha channel of RGBA explanations.
@@
A.3.1 FunnyBirds - ResNet50
Figure A.5: FunnyBirds [12] result of main metrics Completeness (Com.), Background Independence (B.I.), Accuracy, Contrastivity (Con.), Correctness (Cor.).
Method A BI CSDC PC DC D SD TS Com. Cor. Con.
ProtoPNet [8] 0.94 1.00 0.93 0.91 0.92 0.58 0.24 0.46 0.75 0.24 0.46
BagNet [6] 1.00 1.00 0.95 0.98 0.91 0.91 0.76 0.99 0.93 0.76 0.99
B-cos [5] 0.96 0.87 0.93 0.88 0.94 0.86 0.69 0.89 0.89 0.69 0.89
SIC (ours) 0.99 0.99 0.96 0.98 0.96 0.97 0.72 0.95 0.97 0.72 0.95
Table A.4: Results on the FunnyBirds Framework with a ResNet50 backbone. Note that Completeness (Com.) is the mean across Controlled Synthetic Data Check (CSDC), Preservation Check (PC), Deletion Check (DC), and Distractability (D); Correctness (Cor.) equals Single Deletion (SD), and Contrastivity (Con.) equals Target Sensitivity (TS).
(a) Prediction of the representative FunnyBirds [12] test sample. The primary location of relevance is the tail of the bird, while the wings, beak, eyes, and legs show a lower contribution. We observe a variety of poses of representative support samples, e.g. a strong focus on the tail and wings in the first support sample, a focus on the legs, eyes, and wings in the second support sample, and a focus on all visible concepts in the last support sample.
(b) Prediction of another test sample. Similar to (a), all concepts are relevant for the decision making. While the tail appears to be the most important concept for the classification in (a), the eyes seem more important in the specific example. Additionally, the long beak contributes over its whole length, suggesting that the model is sensitive to the appearance of the concept rather than encoding color only.
Figure A.6: Comparison of Explanations.
@@ A.4.4 RSNA - DenseNet121
Figure A.15: Prototypes of SIC trained with a DenseNet121 backbone on RSNA. Bounding boxes indicate where medical staff found occlusion of the lungs. The first three columns are the support samples of controls, the next three are support samples with occlusion present.
@@
A.4.5 VOC - DenseNet121: Prototypes and B-cos Similarities
Figure A.17: Global explanations for SIC trained on Pascal VOC with a DenseNet121 backbone. The respective second rows show the raw contribution maps $\phi^{L}_{j=c}$, in which red denotes positive contributions and blue (absent) negative contributions. The respective third columns present RGBA explanations.
Figure A.18: Heatmap of Intra-class and Inter-class B-cos Similarities for SIC trained on Pascal VOC with a DenseNet121 backbone.